OpenStudy (anonymous):

4. Suppose you administered an anxiety test to a large sample of people and obtained normally distributed scores with a mean of 45 and standard deviation of 4. If Andrew scored 45 on this test. What is his Z score?

OpenStudy (campbell_st):

his z score is 0. \[z = \frac{X - \mu}{\sigma}\] in your question X = 45 mean = 45 SD = 4 so \[z = \frac{45 -45}{4} = 0\]